OpenStudy (anonymous):

Hook’s law.the distance d when a spring is stretched by a hanging object varies directly as the weight w of the object.if the distance is 33 cm when the weight is 3kg what is the distance when the weight is 6 kg

OpenStudy (anonymous):

uhhhh, 66cm??

OpenStudy (shubhamsrg):

well w = kx so w/x = constant => w1/x1 = w2/x2 just plug in values

OpenStudy (anonymous):

for the mass: W=mg for the spring: F=kd the weight and the force of the spring are opposite and the mass is in equilibrium so: mg=kd or W=kd so W and d are proportional (if k is constant). so when W gets twice as big d also gets twice as big. so 66 cm.